MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2006/08/09
OBSERVER: Ohnishi Motomura
CONDITION: Fine, fullmoon
COMMENT: The rotater censer is fixed. The operation restart today! However B3216-B3230 is 3 degree anti-clockwise rotation, and B3231-B3235 is 1.6 degree clockwise rotation due to miss of offset. After B3236 is OK (however same shift renaine. We will reset the offset as soon as possible). 60 sec Dark D1262-1264 and 150sec Dark D1265 are OK.
Compressor gas pressure: MPa
Expected Observation Time: START 18:53 END 06:53


TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB3216-32604526816080
LMC565-5662284200
SMC161-1633193000
ALERT-000
GRB-000
TOTAL---5031523280 (06h28m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
99994599945999
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
99999999880


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
222222222222
Field13Field14Field15Field16
2200


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
322222222000


SPECIAL TARGET NAME

CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
18:00waitingfine
19:00observingfine
20:00observingfine
21:00observingfine
22:00observingfine
22:00observingfine
23:00observingfine
0:00observingfine
01:00observingfine
02:00observingfine
03:00observingfine
04:00observingfine
05:00observingfine
06:00observingfine
06:45observingfine, stop due to dawn


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day